OpenStudy (anonymous):

Kerry bought a pair of athletic shoes for $45. She also bought a pair of colorful shoelaces for the shoes for $4. What percent of the total amount that Kerry paid was the price of the shoelaces? Round to the nearest tenth of a percent, if necessary. A. 8.2% B. 8.9% C. 9.8% D. 11.3%

OpenStudy (anonymous):

so can u help me plz

OpenStudy (anonymous):

well since this is a percent problem you would take the total and pt the amount you trying to find over the total so you you spent a total of $49 shoelaces were $ 4 so you would put the 4/49 to find the percent the shoe laces cost out of the total

OpenStudy (anonymous):

and divide

OpenStudy (anonymous):

45+4= 100/49=2.04*4=8.16 which is equal to 8.2

OpenStudy (anonymous):

percent